# A Survey of Cross-Lingual Alignment: Definitions, Methods, Future Understanding Cross-Lingual Alignment—A Survey

### v1

@@ 1 Introduction
As we define it, cross-lingual alignment means that words or sentences with similar semantics: 1) Are more similar in the representation space than words or sentences with dissimilar semantics. This way of looking at alignment can be defined in “weak” or “strong” terms (see § 2). 2) Allow a prediction head trained on a source language to recognise the relevant patterns in the target language. We argue that this is related to the first view but, importantly, is not identical to it. This allows us to discuss the literature in a new way.
These two criteria are not guaranteed to be fulfilled through unsupervised pre-training, motivating various efforts to improve cross-lingual alignment. We survey important papers in this area between 2019 and 2023. These works propose new training objectives, new pre-trained models, contrastive fine-tuning, or post-hoc adjustments of the embedding space. The vast majority of the methods were developed for and applied to multilingual encoder models, chiefly XLM-R and mBERT.
In this paper, we provide a definition of cross-lingual alignment under both views, how it is currently measured (and weaknesses of these measurements), and an important critique of “strong” alignment as stated by the first view (§ 2); a taxonomy of methods to increase alignment taken from a comprehensive survey of the literature (§ 3); a summary of findings and where the field currently stands (§ 4); and finally a discussion of future research in this area given the developing body of work on generative models (§ 5). As we discuss, generative models pose new and interesting challenges, because we will need to trade-off language-neutral with language-specific factors in new ways.
@@ 2.1 Definitions
To see why cross-lingual alignment under this view is hard to achieve, consider the isomorphism assumption as stated for cross-lingual static embeddings (cf. Vulić et al., 2020). This states that spaces of both languages should have (roughly) the same shape, measured by, for example, relational similarity (Vulić et al., 2020) or eigenvector similarity (Søgaard et al., 2018).
@@
A prediction head trained on a source language should be able to find relevant patterns in the representations of a target language, and classify accordingly. Although it is tempting to think of cross-lingual alignment in terms of simple measures such as cosine similarity, the prediction head works with the full encoder representations as its input, and can (potentially) use subspaces to that effect. Given the many constraints on the representations, it is actually very difficult to attain “cross-lingual alignment”, particularly strong alignment, under View I. However, we can consider language-specific subspaces, as well as features pertaining to specific tasks. For example, some works find subspaces for morphosyntactic aspects (Hewitt and Manning, 2019; Acs et al., 2023), others find directions encoding token frequency (Rajaee and Pilehvar, 2022; Puccetti et al., 2022). Chang et al. (2022) separate types of subspaces by how their means and variances differ in different languages. That is, if both are similar across languages, the axis is language-neutral. If the means differ between languages and/or variances are very different, the axis is language-sensitive.
@@ 2.2 Measuring Cross-Lingual Alignment
is sometimes used (e.g., Libovický et al., 2020) to measure language-specific elements of the representations. In this thinking, if a language classifier trained on the output representations performs poorly, then the model representations are language-neutral. In a sense, this is an even stricter goal than “strong alignment”. However, it neglects that the representations can have both language-neutral and language-specific areas, and that some language-specific information is necessary.
@@ 3 Taxonomy of Alignment Strategies
| Parallel, word-level | Cao et al. (2020); Wu and Dredze (2020); Joint-Align + Norm (Zhao et al., 2021); VECO (Luo et al., 2021); WEAM (Yang et al., 2021); XLM-Align (Chi et al., 2021b); WAD-X (Ahmat et al., 2023) | |
@@
xTune (Zheng et al., 2021); FILTER (teacher model) (Fang et al., 2021); XeroAlign (Gritta and Iacobacci, 2021); Cross-Aligner (Gritta et al., 2022); X-MIXUP (Yang et al., 2022) | FILTER (student model) |
@@
We report on methods for improving zero-shot transfer and increasing cross-lingual alignment. Table 1 shows all included papers, organised by initialisation and data requirements of their proposed objectives. In this section, we discuss each category with examples. We additionally describe categories which are not listed in the table since they would overlap with multiple table cells. We leave out some methods that are less relevant to the overall analysis, though we explain them in Appendix A for completeness.
@@ 3.1 Objectives using Parallel Data
Among pre-trained models, LASER (Artetxe and Schwenk, 2019) is a 5-layer BiLSTM trained on machine translation, with the decoder being discarded. Its successor LASER3 (Heffernan et al., 2022) is a 12-layer Transformer model, but trained using a student-teacher setting, where the teacher is similar to the original LASER. This follow-up also emphasises support for lower-resource languages, training a student for each group of similar languages.
@@ 4 What We Do and Don’t Know
In terms of View I, ‘strong’ alignment is often seen as desirable, but we question whether this should be a main goal. As discussed in § 2, there is a risk of trading off language-specific information. As for downstream tasks, most cross-lingual retrieval tasks query only the target language space, ignoring the language component. Thus, strong alignment may not be necessary for these tasks. LAReQA (Roy et al., 2020) does test for strong alignment specifically, but it is the exception rather than the rule. In tasks with a prediction head, strong alignment would ensure that the condition of View II is met. Indeed, Gaschi et al. (2023) show that strong alignment correlates with better performance on downstream tasks, though they specifically look at classification tasks, no retrieval tasks. However, this does not tell us if strong alignment is necessary for improved transfer performance on these tasks, only that it is correlated.
@@ 5 Multilingual Generative Models
Recently, the field has turned much attention to generative Large Language Models (LLMs). In this space, there are still fewer intentionally multilingual models (e.g., Workshop et al., 2023; Lin et al., 2022), and unfortunately they skew more heavily towards English data than models in our survey. However, we believe that multilingual generation will become increasingly relevant as applications scale. Thus, we point out several areas of future research, including how cross-lingual alignment will interact with multi- and cross-lingual generation.
5.1 Until Now: Encoder-Only Models
So far we have surveyed primarily encoder-only models, though we included a few encoder-decoder models. Encoder-only models transform the inputs into a latent space representation which is typically used by a downstream task “head”. For any tasks where the set of outputs does not depend on the language, the model needs to primarily rely on language-neutral axes of the representations. Intuitively, strong cross-lingual alignment can be helpful here, including more “radical” methods such as mean-centering language-specific embeddings. We have seen this born out by the proliferation of alignment methods and their successes.
5.2 Generative Models
Encoder models are able to predict the most likely tokens to fill a masked position, but encoder-decoder and decoder-only models are more suited to generative tasks due to their architecture.
In the encoder-decoder framework, the encoder outputs a latent space representation, while the decoder predicts the next tokens one-by-one. It is still desirable for the encoder to align the semantics of different languages. Therefore, if we want to use established cross-lingual alignment techniques for these models, the encoder is a natural target for these—while the decoder must focus at least partially on language-specific information in order to generate tokens in the correct language.
Many recent LLMs (Brown et al., 2020; Touvron et al., 2023) are decoder-only models, meaning they use the decoder architecture throughout the model and have no separate encoder layers. Thus, there is no one obvious point at which cross-lingual alignment should be greatest. Further, the classic zero-shot cross-lingual transfer paradigm encounters issues in generative settings, since well-aligned representations can lead to generation in the wrong language (Xue et al., 2021; Li and Murray, 2023).
@@ 5.3 Cross-Linguality in Generative Models
Some works have started to explore cross-lingual alignment in the generative context. For instance, Zhang et al. (2023) tune their model with standard instruction data and interactive translation examples in order to improve both translation and cross-lingual instruction following. Li et al. (2023) tune the hidden representations of the first layer using translated data and contrastive learning, which they combine with cross-lingual instruction tuning. Li and Murray (2023) use two annotated source languages for cross-lingual generation. Other methods, such as Tanwar et al. (2023); Huang et al. (2023); Zhang et al. (2024), rather focus on prompting strategies for cross-lingual transfer.
@@
However, there is plenty of potential for future work in this space. Importantly, training schemes will need to enable the model to focus on language-neutral and language-specific information at the relevant times. This can be combined with sparse fine-tuning approaches such as Hu et al. (2021a), or ideally integrated by those pre-training new multilingual models. The goal is to enable better transfer of information between languages while outputting text in the relevant language.
5.4 Evaluation of Generative Models
@@
There are also some efforts to benchmark multilingual generation (Asai et al., 2023; Ahuja et al., 2023; Gehrmann et al., 2022), but this presents unique challenges compared to both multilingual classification and monolingual generation tasks. These benchmarks include a number of reformulated classification tasks in addition to some translation or summarisation tasks. Classification tasks are not the main strength of generative models, and fine-tuned encoder models may do better there (Lin et al., 2022). Translation and summarisation tasks can be evaluated relatively easily using a reference text, while more open-ended generation is harder to evaluate. Increasingly, papers use ChatGPT as a judge on open-ended generation (e.g., Liu et al., 2023), but this approach has reproducibility issues, and is less likely to work well for low-resource languages.
@@ 6 Conclusions
We have surveyed the literature around cross-lingual alignment, providing a taxonomy of methods. We clarified two main views of the concept, noting how simplistic popular measurements are. We summarised insights from the surveyed methods and related analysis papers. Going forward, new challenges present themselves with respect to multilingual generative models: Simply maximising cross-lingual alignment can lead to wrong-language generation. We thus call for methods that effectively trade-off cross-lingual semantic information with language-specific axes, allowing models to generate fluent and relevant content in many languages.
@@ Acknowledgements
This publication was supported by LMUexcellent, funded by the Federal Ministry of Education and Research (BMBF) and the Free State of Bavaria under the Excellence Strategy of the Federal Government and the Länder; and by the German Research Foundation (DFG; grant FR 2829/4-1). The work at CUNI was supported by Charles University project PRIMUS/23/SCI/023, and by the European Commission via its Horizon research and innovation programme (No. 870930 and 101070350).
@@ A.3 More Tuning with Task Data
DuEAM (Goswami et al., 2021) uses data from the XNLI dataset while targeting semantic textual similarity and bitext mining tasks. The objectives used are Word Mover’s Distance and a translation mining loss. The model performs reasonably well but does not reach the performance of S-BERT.
@@ B.1 What works well?
First, whether the model is newly-trained or modified from a pre-trained model does not appear to determine performance. WordOT, a modified mBERT with an optimal transport objective, yields the best result in its size band. In the next size group, the newly-trained mDeBERTaV3 performs best. This is a model trained with only monolingual data, with the ELECTRA pre-training objective and additional improvement in the form of gradient-disentangled embeddings. XLM-E, which does not have this additional element, does markedly worse than mDeBERTaV3. Only few points behind, ERNIE-M, HiCTL and the JointAlign+Norm method sit at a near-identical performance. All modify an existing model in different ways: InfoXLM uses information theory, ERNIE-M focuses on aligning the attention parameters, whereas JointAlign+Norm looks at the output vector space. In the next group, xTune’s consistency regulation proves highly effective, with ERNIE-M and InfoXLM just behind.
In both zero-shot transfer and translate-train, once we cross the threshold of 1B parameters, XY-LENT is the best available method—we do not know, at this point, if this model would be outperformed by another method being scaled up. Trained from scratch, XY-LENT specifically uses a lot of parallel data that is not only English-centric, which seems to work well. XLM-R lags behind XY-LENT and XLM-E while outperforming its own XL counterpart. Interestingly, mT5, which underperforms in smaller configurations, is competitive in XL size and does very well in XXL.
In the translate-train setting, mDeBERTaV3 again wins its size group. However, in the next larger group of models, X-MIXUP proves the most effective. This method also improves mBERT’s performance by a large margin. X-MIXUP directly addresses representation discrepancies between different languages by linear interpolation between the hidden states of translation pairs. HiCTL, VECO, and ERNIE-M come close to the performance of X-MIXUP on this task, while needing more resources. The contrastive learning approaches in these tables do well (HiCTL, InfoXLM), although they are not necessarily the most performant. We must add the caveat that not all relevant models are listed in the tables, since not all papers report the full XNLI results.

### v2

# Understanding Cross-Lingual Alignment—A Survey

@@ 1 Introduction
As we define it, cross-lingual alignment means that words or sentences with similar semantics: 1) Are more similar in the representation space than words or sentences with dissimilar semantics. This way of looking at alignment can be defined in “weak” or “strong” terms (see § 2). 2) Allow a prediction head trained on a source language to recognise the relevant patterns in the target language. We argue that this is related to the first view but, importantly, is more nuanced. This allows us to discuss the literature in a new way.
These two criteria are not guaranteed to be fulfilled through unsupervised pre-training, thus motivating various efforts to improve cross-lingual alignment. We survey important papers in this area between 2019 and 2023. These works propose new training objectives, new pre-trained models, contrastive fine-tuning, or post-hoc adjustments of the embedding space. The vast majority of the methods were developed for and applied to multilingual encoder models, chiefly XLM-R and mBERT.
In this paper, we provide
a definition of cross-lingual alignment under both views, how it is typically measured—and weaknesses of these measurements—, and an important critique of “strong” alignment as stated by the first view (§ 2);
a taxonomy of methods to increase alignment taken from a comprehensive survey of the literature (§ 3);
a summary of findings and where the field currently stands (§ 4);
and finally a discussion of ongoing and future research in this area as the field focuses more on generative models (§ 5). As we discuss, generative models pose new challenges for cross-lingual alignment, since we need to trade-off language-neutral with language-specific elements in new ways.
@@ 2.1 Definitions
To see why cross-lingual alignment, particularly “strong” alignment, under this view is hard to achieve, consider the isomorphism assumption as stated for cross-lingual static embeddings (cf. Vulić et al., 2020). This states that spaces of both languages should have (roughly) the same shape, measured by, for example, relational similarity (Vulić et al., 2020) or eigenvector similarity (Søgaard et al., 2018).
@@
A prediction head trained on a source language should be able to find relevant patterns in the representations of a target language, and classify accordingly. Although it is tempting to think of cross-lingual alignment in terms of simple measures such as cosine similarity, the prediction head works with the full encoder representations as its input, and can (potentially) use subspaces to that effect. Given the many constraints on the representations, it is actually very difficult to attain “cross-lingual alignment”, particularly strong alignment, under View I. However, we can consider language-specific subspaces, as well as features pertaining to specific tasks. For example, some works find subspaces for morphosyntactic aspects (Hewitt and Manning, 2019; Acs et al., 2023), others find directions encoding token frequency (Rajaee and Pilehvar, 2022; Puccetti et al., 2022).
Chang et al. (2022) separate types of subspaces by how their means and variances differ in different languages. That is, if both are similar across languages, the axis is language-neutral. If the means differ between languages and/or variances are very different, the axis is language-sensitive.
@@ 2.2 Measuring Cross-Lingual Alignment
is sometimes used (e.g., Libovický et al., 2020) to measure language-specific elements of the representations. In this thinking, if a language classifier trained on the output representations performs poorly, then the model representations are highly language-neutral. This is actually an even stricter goal than “strong alignment”. However, it neglects that the representations can have both language-neutral and language-specific areas, and that some language-specific information is necessary.
@@ 3 Taxonomy of Alignment Strategies
Parallel, word-level | Cao et al. (2020); Wu and Dredze (2020); Joint-Align + Norm (Zhao et al., 2021); VECO (Luo et al., 2021); WEAM (Yang et al., 2021); XLM-Align (Chi et al., 2021b); WAD-X (Ahmat et al., 2023); Efimov et al. (2023) | |
@@
We report on methods for improving zero-shot transfer and increasing cross-lingual alignment. Table 1 shows all included papers, organised by initialisation and data requirements of their proposed objectives. In this section, we discuss each category with examples. Additionally, we describe some strategies which are applied across different data requirements and initialisations. We note how methods fit with the two views introduced in § 2. Some methods are left out here because they are less relevant to the overall analysis, though we explain them in Appendix A for completeness.
@@ 3.1 Objectives using Parallel Data
Among pre-trained models, LASER (Artetxe and Schwenk, 2019) is a 5-layer BiLSTM trained on machine translation, with the decoder being discarded. Its successor LASER3 (Heffernan et al., 2022) is a 12-layer Transformer model, but trained using a student-teacher setting, where the teacher is similar to the original LASER. The follow-up also emphasises lower-resource languages, training a student for each group of similar languages.
@@ 4 What We Do and Don’t Know
In terms of View I, ‘strong’ alignment is often seen as desirable, but we question whether this should be a main goal. As discussed in § 2, there is a risk of trading off language-specific information. As for downstream tasks, most cross-lingual retrieval tasks query only the target language space, ignoring the language component. Thus, strong alignment may not be necessary for these tasks. LAReQA (Roy et al., 2020) does test for strong alignment specifically, but it is the exception rather than the rule. In tasks with a prediction head, strong alignment would ensure that the condition of View II is met. Indeed, Gaschi et al. (2023) show that strong alignment correlates with better performance on downstream tasks, though they specifically look at classification tasks, not retrieval tasks. However, this does not tell us if strong alignment is necessary for improved transfer performance on these tasks, only that it is correlated.
@@ 5 Multilingual Generative Models
Recently, the field has turned much attention to generative Large Language Models (LLMs). We believe multilingual capabilities of generative models will be more and more important as applications scale. Thus, we point out several areas of future research, including how cross-lingual alignment may interact with multi- and cross-lingual generation.
5.1 Multilinguality As An Afterthought?
The flagship models (e.g., Touvron et al., 2023; Jiang et al., 2023; OpenAI et al., 2024) in the generative space tend to be English-centric, with some multilingual capability. However, the proportion of multilingual data is often relatively small and a performance drop-off for mid- and low-resource languages is noticeable (e.g., Ahia et al., 2023).
Due to the significant costs of pre-training, only few models have been trained to be intentionally multilingual from the start (e.g., BLOOM, Workshop et al., 2023; XGLM, Lin et al., 2022). Unfortunately, even these skew more heavily towards English data than, e.g., XLM-R, and they tend to underperform compared to, e.g., newer GPT, Llama or Mistral models. mGPT (Shliazhko et al., 2024) has a more balanced language distribution—the model is comparable with XGLM on few-shot classification tasks, but the evaluation on generative tasks is somewhat limited.
Given this research landscape, we see room to 1) explore why efforts like XGLM or BLOOM have been less effective than retrainings of English-centric models, beyond the oft-cited ‘Curse of Multilinguality’ (Chang et al., 2023); and 2) be even more intentional about balancing training data, and including multiple languages. This will necessarily include further data collection efforts as well as additional work on effective multilingual pretraining.
@@ 5.2 Cross-Linguality in Generative Models
However, numerous efforts have recently extended open models to new languages. Examples include EEVE (Kim et al., 2024) for Korean, PolyLM (Wei et al., 2023) with a focus on Chinese, Andersland (2024) for Amharic, or the ongoing Occiglot project (Team, 2024) for European languages. Some models, like Tower (Alves et al., 2024) and ALMA (Xu et al., 2024) are trained with a focus on machine translation. Yet other methods, such as Tanwar et al. (2023); Huang et al. (2023); Zhang et al. (2024), rather focus on prompting strategies to achieve good cross-lingual transfer without necessarily retraining the models.
The Aya Initiative collected instruction data for over 100 languages and varieties, releasing Aya-101 (Üstün et al., 2024), initialised from mT5, and Aya-23 (Aryabumi et al., 2024), initialised from Cohere’s Command R model. The latter paper appeals to the ‘Curse of Multilinguality’ for Aya-101 and other massively multilingual models, and focuses on training Aya-23 with those languages where a larger amount of data is available. By contrast, MaLa-500 (Lin et al., 2024) continues training from Llama-2 and emphasises training on a very broad set of languages.
@@
Zhang et al. (2023) tune their model with instruction data and interactive translation examples in order to improve both translation and cross-lingual instruction following. Li et al. (2023) tune the hidden representations of the first layer using translated data and contrastive learning, which they combine with cross-lingual instruction tuning—this is one of few works on generative models explicitly using a notion of cross-lingual alignment. Li and Murray (2023) use two annotated source languages for cross-lingual generation.
@@
Still, there is plenty of potential for future work in this space. Cross-lingual alignment of representations has not been explored extensively for generative models. Importantly, training schemes will need to enable the model to focus on language-neutral and language-specific information at the relevant times. Training for cross-lingual alignment may be combined with sparse fine-tuning approaches such as LoRA (Hu et al., 2021a). Ideally it should also be integrated by those pre-training new multilingual models. The goal is to enable better transfer of information between languages while outputting text in the relevant language.
5.3 Where to Align a Generative Model?
Cross-lingual alignment as we discussed it has been researched primarily in encoder-only models, and some encoder-decoder models. Encoder-only models transform the inputs into a latent space representation which is typically used by a shallow downstream task “head”. For any tasks where the set of outputs does not depend on the language, the model needs to primarily rely on language-neutral axes of the representations. This is clearly the case for the classification and sequence labelling tasks where encoder models shine. The task head has somewhat limited opportunity to transform the encoder representations. Intuitively, strong cross-lingual alignment can be helpful here, including more “radical” methods such as mean-centering language-specific embeddings. We have seen this born out by the proliferation and success of cross-lingual alignment methods in encoders.
In an encoder-decoder framework, there is an explicit delineation with different architectures, where the encoder still outputs a latent space representation, while the decoder predicts the next tokens one-by-one. Cross-lingual alignment techniques from an encoder context can be naturally applied to the encoder of an encoder-decoder model. However, there could be risks—“radical” alignment methods, e.g., mean-centering language-specific embeddings from the encoder, might suppress necessary language-specific information for the decoder, or “break the link” of the encoder to the decoder. Nevertheless, this is certainly a worthwile avenue to explore.
In decoder-only models, there is no such architectural delineation. Thus, there is no one obvious point at which cross-lingual alignment should be greatest. Further, the classic zero-shot cross-lingual transfer paradigm may encounter issues in generative settings, since strongly-aligned representations can lead to generation in the wrong language (Xue et al., 2021; Li and Murray, 2023).
That said, interpretability studies suggest that even in decoder-only models, early layers may perform a kind of encoding role. For instance, Wendler et al. (2024) use an early exiting technique known as logit lens (nostalgebraist, 2020) to illustrate at which layers in the model translating logits to tokens starts to produce reasonable outputs. They demonstrate that Llama-2 “works” in English until late in the model, and only then surfaces tokens from the target language. Meanwhile, Chuang et al. (2024) work only in English but detect hallucinated vs. correct outputs using ‘differential likelihoods’ of early-exited tokens throughout the model. Both papers suggest inflection points in mid-to-late layers, where the correct output language or token starts to become more likely. If we can find a kind of ‘implicit encoder’ in this way, this could become a target for cross-lingual alignment methods.
5.4 Evaluation of Multilingual Generation
Benchmarking multilingual generation presents unique challenges compared to both multilingual classification and monolingual generation tasks.
@@ 5.4 Evaluation of Multilingual Generation
Several benchmarks have been proposed in the past (Asai et al., 2023; Ahuja et al., 2023; Gehrmann et al., 2022). However, these benchmarks include a number of reformulated classification tasks in addition to some translation or summarisation tasks. Classification tasks are not the main strength of generative models, and fine-tuned encoder models often do better there (Lin et al., 2022). Translation and summarisation tasks can be evaluated relatively easily using a reference text—which is likely why they were included in the benchmarks.
However, due to the open-ended nature of many generation tasks, even in high-resource languages they are non-trivial to evaluate. Increasingly, researchers use ChatGPT as a judge on open-ended generation (e.g., Liu et al., 2023), but this approach may not be reproducible due to frequent changes of the ChatGPT model, and it is less likely to work well for low-resource languages. Note also that care should be taken to avoid leaking evaluation data to proprietary models (Balloccu et al., 2024).
@@ 6 Conclusions
We have surveyed the literature around cross-lingual alignment, providing a taxonomy of methods. We clarified two main views of the concept, noting how simplistic popular measurements are. We summarised insights from the surveyed methods and related analysis papers. Going forward, new challenges present themselves with respect to multilingual generative models: Simply maximising cross-lingual alignment can lead to wrong-language generation. We thus call for methods which effectively trade-off cross-lingual semantic information with language-specific axes, allowing models to generate fluent and relevant content in many languages.
@@ Acknowledgements
This publication was supported by LMUexcellent, funded by the Federal Ministry of Education and Research (BMBF) and the Free State of Bavaria under the Excellence Strategy of the Federal Government and the Länder; and by the German Research Foundation (DFG; grant FR 2829/4-1). The work at CUNI was supported by Charles University project PRIMUS/23/SCI/023.
@@ A.3 More Tuning with Task Data
DuEAM (Goswami et al., 2021) uses data from the XNLI dataset while targeting semantic textual similarity and bitext mining tasks. The objectives used are Word Mover’s Distance and a translation mining loss. The model performs reasonably well but does not reach the performance of S-BERT. Efimov et al. (2023) do not tune directly with task data, but they do combine a Cao et al. (2020)-style alignment objective with task fine-tuning in a multi-task-learning objective, with mixed success.
@@ B.1 What works well?
First, whether the model is newly-trained or modified from a pre-trained model does not appear to determine performance. WordOT, a modified mBERT with an optimal transport objective, yields the best result in its size band. In the next size group, the newly-trained mDeBERTaV3 performs best. This is a model trained with only monolingual data, with the ELECTRA pre-training objective and additional improvement in the form of gradient-disentangled embeddings. XLM-E, which does not have this additional element, does markedly worse than mDeBERTaV3. Only few points behind, ERNIE-M, HiCTL and the JointAlign+Norm method sit at a near-identical performance. All modify an existing model in different ways: InfoXLM uses information theory, ERNIE-M focuses on aligning the attention parameters, whereas JointAlign+Norm looks at the output vector space. In the next group, xTune’s consistency regulation proves highly effective, with ERNIE-Mlarge and InfoXLM just behind.
In both zero-shot transfer and translate-train, once we cross the threshold of 1B parameters, XY-LENTXL is the best available method—we do not know, at this point, if this model would be outperformed by another method being scaled up. Trained from scratch, XY-LENT specifically uses a lot of parallel data that is not only English-centric, which seems to work well. XLM-RXXL lags behind XY-LENTXL and XLM-EXL while outperforming its own XL counterpart. Interestingly, mT5, which underperforms in smaller configurations, is competitive in XL size and does very well in XXL.
In the translate-train setting, mDeBERTaV3 again wins its size group. However, in the next larger group of models, X-MIXUP proves the most effective. This method also improves mBERT’s performance by a large margin. X-MIXUP directly addresses representation discrepancies between different languages by linear interpolation between the hidden states of translation pairs. HiCTL, VECO, and ERNIE-Mlarge come close to the performance of X-MIXUP on this task, while needing more resources. The contrastive learning approaches in these tables do well (HiCTL, InfoXLM), although they are not necessarily the most performant. We must add the caveat that not all relevant models are listed in the tables, since not all papers report the full XNLI results.